A mother is 33 years older than her twin children. In 3 years her age will be twice the combined ages of the twins. How old are the twins now?
Answer by rothauserc(4718)

let x be age of her twins and y be the age of their mother
:
1) y = x + 33
2) y + 3 = 2(x + x +3)
:
using equation 1 substitute for y in equation 2
:
x+33 +3 = 2(2x+3)
:
x + 36 = 4x + 6
:
3x = 30
:
x = 10
:
y = 10 + 33 = 43
:
****************************
the twins are 10 years old
their mother is 43 years old
****************************
: